Instrument Announcements

A tag for the instrument must also be used.

GNIRS Temporarily Limited to Cross Dispersed Mode

April 25th, 2022

In January 2022, itermittent problems began to occur with GNIRS' filter wheel two (FW2). This was found to be due to a faulty ball bearing, which increased drag and friction on the wheel assembly. In order to prevent damage to FW2 or other components internal to GNIRS, a decision was made to avoid FW2 movements entirely. As a result, as of March 15th, 2022, observations were limited to cross dispersed mode only. The faulty ball bearing is to be substituted when GNIRS is warmed up to install the IFU at the end of April 2022. 

April 2022 GHOST Update

April 1st, 2022

The GHOST project commissioning has begun, with the Canadian and Australian teams able to travel to the summit following the easing of COVID-19 related restrictions. Optical bench assembly has begun, with alignment tests scheduled for late April. Twilight tests are to take place in mid-May. On-sky science commissioning is planned to take place in mid-late June, 2022. 

Webpage on GMOS-N detector status

March 18th, 2022

A new webpage with information on the current status of the GMOS-N detector array is now available here. The webpage provides details on the recent hot column features (see March 10 announcement) and updated recommendations for dither strategies to minimize the impact of the hot columns on science data. Any future updates from the ongoing characterization of the hot columns will be posted on this webpage. 

GMOS-N status update

March 10th, 2022

The GMOS-N detector issue reported on March 2 has largely been resolved by the full thermal cycle performed between March 2 and 3. The extended bright columns on amplifiers 7 and 12 are no longer present. The narrow bad column of hot pixels on amplifier 5 persists and saturates in longer science exposures. The following figure shows the hot column in an overscan-subtracted bias image (as marked by the blue arrow). 

Bias image showing the new hot column

New GMOS-N bias features

March 2nd, 2022

The GMOS-N detector has shown new bias features since an uncontrolled warm-up of the detector on February 26, 2022 . These features consist of a narrow bright column on amplifier 5, and broader bright columns on amplifiers 7 and 12. As a first measure to address this new issue, a full thermal cycle of the detector is being performed starting from today. GMOS will be unavailable for the next couple of nights until the thermal cycle is complete.

F-2 is back in operations

January 29th, 2022

F-2 was successfully checked on January 26th after an intevention that lasted for most of January. New JH and HK gratings were installed and its performance will be assessed in the following days. Additionally, the  J_low filter is now permanently installed.

F-2 planned shutdown

November 19th, 2021

An intervention of Flamingos-2 is planned between January 3rd and 21st where the instrument will be offline. The following tasks are planned:

MOS wheel check and preventive work.
Replace the JH/HK spectroscopic filters with higher throughput ones.
Modify the Lyot wheel to allow the installation of the J-low and Y filters instead of the Hartmann masks.
Increase the number of layers in the dewars insulation, to allow more thermal stability.